OpenStudy (anonymous):

Burning Butane C4H10,from a Camping Gaz Container raised the temperature of 200g water from 18 (degrees celcius) to 28 (degree celcius).The Gaz container was weighed before and the loss in mass was 0.29g.Estimate molar enthalpy change of Combustion of Butane

OpenStudy (chmvijay):

Q=M *C *T =0.2kg *4.28* (28-18) Q=8.56 so 0.29 gram of combustion of butane given 8.56 KJ of energy. convert 0.29 gram butanol to moles to get molar enthalapy change moles=gram/mol wt= 0.29/58.12 =0.004989moles 0.004989 of butane = 8.56 KJ 1mole of butane=8.56*1/0.004989=1715.7KJ/mol

OpenStudy (chmvijay):

where c is constant specific heat of water and weight loss indicates the amount of butane reacted!!!

OpenStudy (chmvijay):

if u have any doubt u can ask!!!1

OpenStudy (anonymous):

thank you thats a bit close but the mark scheme`s answer is 1680kj/mol.i wonder how?

OpenStudy (chmvijay):

yaaa thank you !!!

OpenStudy (anonymous):

i calculated and i got 1676 kJ/mol, but chmvijay probably rounded some numbers as did i so we didnt get 1680 kJ/mol

OpenStudy (chmvijay):

yaaa true sir!!!

OpenStudy (anonymous):

Yes true :)